solve the equation3(x-1)-2(2x)=-6
Answer by piglet162431(90)

Simplify the distributive property first.
3x-3-4x=-6
Now, plus 3 to each side
3x-4x=-3
-1x=-3
x=3
Check.
9-3-12=-6
6-12=6
-6=-6